# references.bib {refsection}

### v3

@@ 2.2.3 Data Analysis
The research team proceeded through several steps during the data analysis that included familiarization with the data, open coding, and axial coding. First, each member of the research team read 10 interviews verbatim to gain familiarity with the participants’ experiences and perceptions. Then, each member of the research team engaged in open coding, breaking the data down into separate elements or units, and examined line by line to identify initial concepts Strauss1990. After coding all transcripts and agreeing on first-order codes, the researchers reviewed and reevaluated their work to guarantee accurate coding. During the analysis of the data, the researchers engaged in a process of comparing and contrasting their interpretations until a consensus was reached, ensuring that each member of the team was satisfied with the categorization. These concepts were then compared for similarities and differences, and grouped into categories. In the second phase, each member then engaged in axial coding with the same 10 transcripts and the categories that emerged from open coding were further developed and refined by exploring the relationships between them. Open coding categories were organized into related clusters, and connections were established among them during this phase of the analysis. Again, the research team discussed all axial codes and higher-order themes until consensus was achieved. Lastly, each member of the research team swapped 10 transcripts that were different from the original 10 we analyzed, and reviewed the data and subsequent open and axial codes to ensure trustworthiness and reliability. During this process, we flagged codes and themes that we needed more discussion on due to conflicting interpretations. In the final phase, the second, third, and fifth-listed authors finalized codes and conducted a category audit to ensure standards of trustworthiness and reliability were met. Through this iterative process of open and axial coding, the underlying structure of the data was uncovered, and five higher order themes and subthemes emerged from the data.
3 Results
3.1 General population quantitative survey results
Agreement with provided definitions:
We found that the Gen Pop. respondents generally agreed (81-96%) with the standard definitions provided for AI, Machine Learning, Fairness and bias (Fig. 2). Please see Appendix A.1.1 for the specific definitions provided. Of the provided definitions, bias in AI had the highest levels of strongly disagree, disagree and neutral.
Figure 2: General population responses to likert scale questions on agreement with AI definitions, familiarity with AI, trust, benefits, concerns, country-level disparities, and colonialism. a) Responses to questions poised in a likert scale format b) Responses to questions poised in a semantic differential scales format.
Familiarity with AI and use cases:
Most of the Gen Pop. respondents were familiar with AI and its associated use cases with 11% indicating they were familiar with AI and use it regularly, 43% indicating that they had read a lot about AI and used it themselves, 31% indicating they were moderately familiar with AI but have not read much about it or used it themselves, 14% indicating they had heard of AI conceptually but had not used it, and only 2% indicating that they were not familiar at all (Fig.2b). We purposefully screened for respondents who had some familiarity of AI and could therefore provide knowledgeable, informed answers/opinions to the follow on questions. However this also biases our results against the section of the population which may be less educated and less aware of AI.
Trust, benefits and concerns on AI/ML tools:
@@ 3.1 General population quantitative survey results
Gen Pop. respondents had high levels of trust for both AI and ML definitions with most indicating they were likely or extremely likely to trust a health intervention if it used AI/ML (74-78%). Seventy-two percent of the Gen Pop. believed that AI was likely or extremely likely to be a force for health empowerment in their countries and about 98% indicated that AI for health had benefits for people in their country. However, while acknowledging the potential benefits of AI, 89% also acknowledged concerns around AI use for health with 47% indicating that they were moderately, very or extremely concerned around the use of AI for health. When asked whether they would use an AI tool that had been shown to perform less accurately for a group of people, only 33% indicated they would still use it, with 35% indicating they might use it depending, and 31% saying they would not use it at all (Fig.2a).
Benefits to country, Inter-country disparities and colonialism:
@@
Case studies around data collection projects for development of AI tools:
@@
Figure 3: Participant responses to case studies around data collection projects for development of AI tools. Respondents were randomly assigned to rate HealthMax, an international organization (a,b) or HealthMax Afrique, a local organization (c,d), along positive axes (a,c) and negative axes (b,d)
Case study around perceptions of AI chatbots for health
We provided respondents with a case study on LLM interactions. Specifically, respondents were shown a question by a “user” posed towards an LLM, and the response by an LLM model. Respondents were shown a base question describing the user’s symptoms and request for advice about what to do next and the LLM response to the question. They were asked to assume they were the user and respond to useability and adherence to the LLM recommendations. Note that the Gen Pop. participants are not considered experts and obtaining their responses on whether or not the LLM answered the question and level of accuracy was primarily to gauge the perceptions users had towards the LLM generated response. These responses do not provide an indication of how accurate the LLM was. Figure 4 shows the results as a percentage with 61% of respondents indicating the response was fully accurate, and 48% indicating it fully answered the question (Fig. 4a). However, a significant portion of respondents (52%) indicated it only partially answered the question (48%) or did not answer the question at all (4%), while 38% indicated the response was only partially accurate and 1% indicated it was not accurate. Most respondents (79%) indicated that the diagnosis was in the list of suggestions. On usability, most respondents (64%) described the length of the response as just right, while 33% indicated it was too long. Only 3% felt the response was too short (Fig. 4a).
Figure 4: Gen Pop. participant responses to chatbot health recommendations. a) Respondents rating LLM responses for answer level, accuracy, length of information, and whether or not the diagnosis is in the list of suggestions. b) Rating of helpfulness of the added local augmentations and religion and gender augmentations, c) Adherence level to LLM-provided recommendations as next steps and how they change with different augmentations. Note that the Gen Pop. population are not considered experts and their responses on whether or not the LLM answered the question and level of accuracy was primarily to gauge the perceptions users had towards the LLM-generated response and not to provide indication of how accurate the LLM was. Note: the last 4 options in the base recommendations of (c) did not appear in the base follow up options. The last option in the location augmentation did not appear in the location followup options.
Respondents were then provided augmentations to the questions that included the location of the user and the change in LLM response based on the location (Fig. 4b). Most Gen Pop. respondents (54%) felt that the response including location context was helpful, 38% indicated it was somewhat helpful and 8% indicated it was not helpful. Respondents were then provided augmentations that added gender and religion of the user to the question and the corresponding LLM contextual response (Fig. 4b). Sixteen per cent indicated this addition was not helpful, 39% indicated it was somewhat helpful or not helpful, and 45% indicated that it was very helpful. We looked at adherence to the LLM recommendations on next steps and found that irrespective of augmentation (none/base, location, gender and religion), most would follow LLM recommendations to see a doctor, nurse, or other health practitioner as next steps (23%-35%), followed by getting plenty of rest (14%-19%), and drinking plenty of fluids (13%-17%) (Fig. 4c). Taking over-the-counter medications came next with the base response, however with local augmentation recommendation, asking neighbors or friends was the next most popular response, and with gender and religious augmentation, seeking spiritual guidance from religious leaders was the next most popular response Fig. 4c. Only about 5-8% indicated they would seek information from elsewhere, indicating high trust in LLM generated recommendations.
Types of questions respondents would ask an AI Chatbot:
Respondents were asked to optionally list 3 examples of questions they would ask an AI chatbot. This generated 520 questions which were analyzed and grouped into categories in a word cloud (Fig. 5). The top categories around which questions fell into were 1) achieving health goals, e.g., “…I want to start a fitness routine. What exercises and workout schedule would be most effective for achieving my specific health goals.” [Rwanda, 30-39, Male]; (2) causes of diseases, e.g., “A sharp pain in the upper left side of the body. What are the possible causes and solutions?” [Ghana, 30-29]; and (3) symptoms of diseases, e.g., “What are the symptoms of a certain medical condition?” [Kenya, 21-29].
Figure 5: Word cloud distribution of question types that Gen Pop. respondents would ask an AI chatbot, with examples of questions submitted. Larger size indicates high occurrence of questions type.
3.2 Expert responses to pre-interview surveys
Expert responses on best and worst data collection practices:
@@ 3.2 Expert responses to pre-interview surveys
Prior to participating in the in-depth interviews, experts were asked to fill out pre-surveys where they listed best and worst data collection practices. Responses are summarized in Table 2
Table 2: Expert best and worst data collection practices
@@
Expert responses to the impact of current barriers to AI for clinical usage:
@@
3.2.1 Comparisons of expert and general population survey on AI axes of disparities

### v4

@@ 3.2.3. Data Analysis
The team met either weekly or bi-weekly for six months to discuss the data process, the insights that were emerging, and engaged in critical discussion to ensure that the analysis codes accurately reflected the data. The research team proceeded through several steps during the data analysis that included familiarization with the data, open coding, and axial coding. Each member of the thematic analysis team engaged in a round of open coding. In the second phase, each member then engaged in axial coding and the categories that emerged from open coding were further developed and refined by exploring the relationships between them. The research team discussed all axial codes and higher-order themes until consensus was achieved. Lastly transcripts were swapped such that each was reviewed by at least 1 other researcher to ensure trustworthiness and reliability. Two of the researchers finalized codes and conducted a category audit to ensure standards of trustworthiness and reliability were met. Through this iterative process five higher order themes and subthemes emerged from the data.
4. Results
4.1. General population quantitative survey results
4.1.1. Trust, benefits and concerns on AI tools:
@@
Gen Pop. respondents had high levels of trust for AI definitions with most indicating they were likely or extremely likely to trust a health intervention if it used AI (74-78%). Seventy-two percent of the Gen Pop. believed that AI was likely or extremely likely to be a force for health empowerment in their countries and about 98% indicated that AI for health had benefits for people in their country. However, while acknowledging the potential benefits of AI, 89% also acknowledged concerns around AI use for health with 47% indicating that they were moderately, very or extremely concerned around the use of AI for health. When asked whether they would use an AI tool that had been shown to perform less accurately for a group of people, only 33% indicated they would still use it, with 35% indicating they might use it depending, and 31% saying they would not use it at all (Fig.2).
Figure 2. General population responses to trust in AI, benefits, concerns, country-level disparities, and colonialism. a) Responses to questions poised in a likert scale format b) Responses to questions poised in a semantic differential scales format.
4.1.2. Benefits to country, Inter-country disparities and colonialism:
@@
4.1.3. Case study around perceptions of AI chatbots for health
We provided respondents with a case study on LLM interactions. Specifically, respondents were shown a question by a “user” posed towards an LLM, and the response by an LLM model. Respondents were shown a base question describing the user’s symptoms and request for advice about what to do next and the LLM response to the question. They were asked to assume they were the user and respond to useability and adherence to the LLM recommendations. Note that the Gen Pop. participants are not considered experts and obtaining their responses on whether or not the LLM answered the question and level of accuracy was primarily to gauge the perceptions users had towards the LLM generated response. These responses do not provide an indication of how accurate the LLM was. Figure 3 shows the results as a percentage with 61% of respondents indicating the response was fully accurate, and 48% indicating it fully answered the question (Fig. 3). However, a significant portion of respondents (52%) indicated it only partially answered the question (48%) or did not answer the question at all (4%), while 38% indicated the response was only partially accurate and 1% indicated it was not accurate. Most respondents (79%) indicated that the diagnosis was in the list of suggestions. On usability, most respondents (64%) described the length of the response as just right, while 33% indicated it was too long. Only 3% felt the response was too short (Fig. 3).
Figure 3. Gen Pop. participant responses to chatbot health recommendations: Respondents rating LLM responses for answer level, accuracy, length of information, and whether or not the diagnosis is in the list of suggestions.
Figure 4. Gen Pop. participant responses to chatbot health recommendations: Rating of helpfulness of the added local augmentations and religion and gender augmentations
Figure 5. Gen Pop. participant responses to chatbot health recommendations: Adherence level to LLM-provided recommendations as next steps and how they change with different augmentations. Note that the Gen Pop. population are not considered experts and their responses on whether or not the LLM answered the question and level of accuracy was primarily to gauge the perceptions users had towards the LLM-generated response and not to provide indication of how accurate the LLM was. Note: the last 4 options in the base recommendations of (c) did not appear in the base follow up options. The last option in the location augmentation did not appear in the location followup options.
Respondents were then provided augmentations to the questions that included the location of the user and the change in LLM response based on the location (Fig. 4). Most Gen Pop. respondents (54%) felt that the response including location context was helpful, 38% indicated it was somewhat helpful and 8% indicated it was not helpful. Respondents were then provided augmentations that added gender and religion of the user to the question and the corresponding LLM contextual response (Fig. 4). Sixteen per cent indicated this addition was not helpful, 39% indicated it was somewhat helpful or not helpful, and 45% indicated that it was very helpful. We looked at adherence to the LLM recommendations on next steps and found that irrespective of augmentation (none/base, location, gender and religion), most would follow LLM recommendations to see a doctor, nurse, or other health practitioner as next steps (23%-35%), followed by getting plenty of rest (14%-19%), and drinking plenty of fluids (13%-17%) (Fig. 5). Taking over-the-counter medications came next with the base response, however with local augmentation recommendation, asking neighbors or friends was the next most popular response, and with gender and religious augmentation, seeking spiritual guidance from religious leaders was the next most popular response Fig. 5. Only about 5-8% indicated they would seek information from elsewhere, indicating high trust in LLM generated recommendations.
4.1.4. Case studies around data collection projects for development of AI tools:
@@
Figure 6. Participant responses to case studies around data collection projects for development of AI tools. Respondents were randomly assigned to rate HealthMax, an international organization (a,b) or HealthMax Afrique, a local organization (c,d), along positive axes (a,c) and negative axes (b,d)
4.2. Expert responses to pre-interview surveys
4.2.1. Expert responses on best and worst data collection practices:
@@
Prior to participating in the in-depth interviews, experts were asked to fill out pre-surveys where they listed best and worst data collection practices. Responses are summarized in Table 2.
Table 2. Expert best and worst data collection practices
@@
4.2.2. Expert responses to the impact of current barriers to AI for clinical usage:
@@
4.3. Findings of expert qualitative interviews
Five major themes emerged from the inductive analysis of expert IDIs namely (1) trust in AI in Africa, (2) opportunities for AI to address health in Africa, (3) intersectional inequities and systemic factors impacting health equity in Africa, (4) social and ethical considerations for developing and deploying AI solutions in health in Africa, and (5) capacity building for effective implementation and adoption of AI solutions in health in Africa. The following section provides a summary of each theme and sub themes with representative quotes. A full writeup of the in-depth thematic analysis can be found in A.2.2.
4.3.1. Trust In AI in Africa
Participants overwhelmingly shared perceptions around sources of mistrust, citing historical marginalization, colonialism, misunderstanding of AI technologies, and a lack of transparency, fairness and accountability as sources. As P25, a Medical Doctor from Nigeria put it, ”a lot of times things that are seen to be coming from the Global North are viewed with a lot of skepticism.” Many participants also highlighted how the legacy of colonialism has fostered a belief that Western innovations, even those aimed at improving healthcare may have ulterior motives, such as continued control, or data exploitation. As P006, a Public Health Researcher in Tanzania stated,…”so it can also be perceived that way, that these people are still trying to continue to control us, tell us what to do, predict how things are going to be in terms of our health and everything”. Through the lens of colonialism and marginalization in Africa, other themes related to misunderstanding of technology and lack of transparency were salient among these experts. As P030 a public health research observed, “There is a lot of misinformation and myths around AI”. These baseline concerns over the use of AI in healthcare are exacerbated by perceptions of biases and lack of accountability with automated systems. In order to tackle these concerns around trust, participants identified the need to build trust for AI use through community-driven approaches, and they proposed methods to foster trust through commitments, education, and good data practices. Most participants stressed the importance of involving local communities to ensure culturally and contextually relevant solutions for AI for health in Africa, which will facilitate adoption of AI. This entails involving communities directly in the problem scoping and development phase, as stated by P059, a Community Health Worker from Kenya ”if the communities are engaged at the point of the system then they can provide background information on the region the communities come from, the kind of cases that have been there, the kind of environment they live in and even the cultural perspectives that affect health”. The data illustrated how essential it is to address concerns about data ownership and sovereignty, ensuring data is used responsibly, and that building trust must include establishing long-term commitments, engaging with stakeholders to showcase the benefits of using AI tools in health, and providing assurance of the effectiveness of these models.
4.3.2. Opportunities for AI to Address Health in Africa
Another major theme that emerged was the expert suggestions of opportunities for AI to address health concerns in Africa, encompassing their views on the desired features of AI solutions for better adoption. As one public health researcher from Nigeria explained, “I think that will be a game changer instead of taking time off these repetitive administrative tasks that the healthcare system [is] currently burdened with and allowing health workers to work at the scope of practice”. Overall, experts highlighted AI’s potential to democratize healthcare, namely the idea that AI could offer greater autonomy to African communities, reducing reliance on outside support by providing tools for self-diagnosis, prevention, and community-based care. P038, a Machine Learning Practitioner/Engineer summarized this sentiment, stating, “We believe that providers and the AI can work really well together. The AI is not looking to replace their expertise, but when they work together, they’re actually augmenting each other”. Participants suggested opportunities for AI to address health concerns in Africa, encompassing their views on the desired features of AI solutions for better adoption. In addition to being contextually relevant, participants emphasized that accuracy and consistency are crucial factors before deploying AI tools in health.
4.3.3. Intersectional Inequities and Systemic Factors Impacting Health Equity in Africa
Experts described various factors, both individual and systemic, that impact health equity in Africa, as well as proposals for addressing existing inequities. Experts broadly identified social determinants of health, infrastructure limitations, and socioeconomic disparities stemming from the history of colonialism as key contributors to health inequities. As P018,a Medical Doctor, noted, “Income is a major factor that contributes to health inequities. People who are poor are less likely to have access to healthcare, and they are less likely to receive quality healthcare when they do have access”. A common theme from the experts centered around the a social-economic stratified system where lower economic areas could greatly benefit from AI in healthcare, but concerns about equitable models in poorer communities was a top of mind concern. These were further compounded by the intersection of other historically oppressed and marginalized identities and belief systems. For example, gender identity was frequently cited as a source of inequity in healthcare access. As P073, a Medical doctor in Nigeria summarized, ”There are some places where the cultural norms do not allow for equitable access to quality healthcare. So for instance, we have some cultures that prohibit a woman . . . to seek healthcare in the health facilities”. The education disparities, in addition to diversity of languages spoken on the continent, also led to further disenfranchisement from groups who do not speak the same language or have access to education, creating additional barriers to healthcare. Some participants also mentioned inequity in healthcare access for sexual minorities, mostly due to stigma and discrimination: “when you actually think about the LGBTQ communities, they actually face a lot of health inequities in general because of stigma and judgment. Those are the key population groups, I think, in Africa that are impacted by the health inequities” - P038 [30-39, USA, Machine Learning Practitioner/Engineer]. Race, ethnicity and disability status also emerged as axis of identity that were negatively impacted by healthcare and several participants expressed hope and healthy skepticism for the role AI could play in mitigating these forms of discrimination. Experts stressed that while targeted interventions can address specific barriers like geographical limitations or lack of funding, there is a need for a comprehensive approach to tackle health inequities in Africa in a sustainable way. The role of governments and private institutions were also seen as critical to addressing health inequities in Africa. Participants, for the most part, pointed to the critical role of government and policy in addressing health inequities in Africa. Many called for better health resource allocation in order to “avail the healthcare service throughout the population on equitable basis” (P008 [30-39, Ethiopia, Public Health Researcher]). Participants also described AI-aided decision making as a solution for improved health outcomes.
4.3.4. Social and Ethical Considerations
Experts articulated the importance of addressing social and ethical factors when introducing AI technologies in healthcare in Africa. One such considerations is the need to develop localized solutions that consider social behaviours and cultural sensitivities. This includes “understanding the context like the social behavior, the context of healthcare in that particular country, the leadership structure as well” (P006 [30-39, Tanzania, Public Health Researcher]) and “making sure the model eliminates all these” (P006 [30-39, Tanzania, Public Health Researcher]). Many participants highlighted the necessity of developing localized solutions that are sensitive to social behaviors, advocating for community engagement throughout the development and deployment of AI solutions. As P038, a Machine Learning Practitioner/Engineer noted, “One of the things is having someone local on the ground. I do believe that if you’re going to have a program that’s deployed in a particular community, you need someone who’s actually familiar with that community involved in the program right from the get go”. Experts also emphasized addressing existing inequities by deploying AI equitably and prioritizing language accessibility. The data underscores that the safe and effective implementation of AI in healthcare in Africa hinges on ethical considerations, bias mitigation strategies, establishing clear regulatory frameworks, and addressing data representativeness, scarcity, collection and privacy. Many participants stressed the need to ensure that AI is deployed in a way that reaches vulnerable populations who are most in need of the technology, as well as taking steps to ensure accessibility in terms of language. P010 [30-39, Kenya, Research and Policy Analyst] noted how AI solution providers need to “really look at having a balance between market profitability and how best to ensure that, other than profits, . . . how does this really get into the general population who need it the most”.
4.3.5. Capacity Building for Effective Implementation and Adoption of AI Solutions in Health in Africa
Experts highlighted strategies to overcome the challenges associated with developing and integrating AI into the healthcare sector in Africa. These strategies encompass enhancing data collection and infrastructure, bridging the digital literacy gap, fostering strong local institutions, and implementing comprehensive training programs and awareness campaigns. This includes promoting open data initiatives and encouraging data sharing among the scientific community as P028 Head of Engineering in South African stated, “I think the open data movement of trying to promote, publishing more data and making that widely available, can go a long way to building better data sets”. Experts indicated that such efforts are crucial for nurturing the African AI ecosystem and building public trust for successful AI adoption in Healthcare in Africa. Investing in local data infrastructure, including “cloud hosting” and data storage units was a common sentiment among participants. For P056, Public Health Researcher, “we need to develop our own databases, our own data storage units to store our own data in Africa, then we’ll be able to use this data to develop algorithms that will . . . benefit the local communities”. Most participants believe that collaboration across stakeholders – including users, local communities, and educational institutions – is vital in driving this development. As P010 a Research and Policy Analyst from Kenya articulated, fostering partnerships and networking across these groups is essential for “helping grow the capacities in Africa, the knowledge”. A USA based Digital Health Lead for an NGO echoed this sentiment, asserting that establishing strong “local institutions and research, whether public or private”, is “a key driver towards…successfully implementing and maintaining AI”
5. Discussion
Our study identified expert and general population perspectives on factors contributing to health inequities in Africa and the potential for AI applications in healthcare to improve population health and reduce health inequities in Africa. A major recurring theme identified relates to the need for contextually-grounded, culturally-aware, and community-driven approaches. For example, this includes the need to understand specific social and structural causes of health inequities, including those related to infrastructure, healthcare access, and local political, regulatory, and governmental environments (Qoseem et al., 2024; Owoyemi et al., 2020; Oladipo et al., 2024); the need for a participatory approach to problem formulation grounded in the needs of the specific communities and contexts in which a system is intended to be used (Martin Jr et al., 2020; Suresh et al., 2024; Passi and Barocas, 2019); as well as contextualized approaches to conceptualizing, measuring, and mitigating issues related to algorithmic fairness, bias, and transparency (Sambasivan et al., 2021; Asiedu et al., 2024). Regarding the potential for AI to promote health equity, participants noted that AI systems and interventions designed to address specific clinical needs (e.g., tools for patient registration, clinical decision support, or diagnostics), structural barriers (e.g., healthcare access), and context-specificity (e.g., augmentation of LLM prompts) may be effective. Furthermore, participants highlighted that, in order to realize this potential, it is critical to address complex structural and social challenges related to infrastructure and capacity building, trust and community attitudes in the context of colonial history, as well as technical limitations related to biases in datasets and AI systems.
Call to Action
: Based on our findings, we present a call to action for AI practitioners, policy makers and researchers in Africa and around the world to consider when deploying AI tools for health in Africa.
  • Increase the use of participatory research at a local level: Participatory research at the local level is the most direct way to develop AI solutions that are grounded in the local contexts and address issues around trust, effective solutions, data representativeness and ethical data collection. Such participatory approach can help bridge the gap between expert concerns and general population optimism thus fostering better adoption.
  • Fund and develop infrastructure for local data centers and compute, given the sensitive nature of health data: Local infrastructure is paramount to address concerns around data sovereignty, ownership, security and privacy as well as improve access to representative datasets for building more unbiased tools.
  • Build capacity to develop home-grown AI for health datasets and tools: this is key to building culturally aware AI solutions that incorporate local knowledge and customs.
  • Apply the expert-sourced recommendations we delineate in Table 2 for data collection practices
  • Improve education and awareness of both AI opportunities and locally-relevant fairness considerations.: this is important to bridge the gap between expert concerns and general population optimism towards AI in healthcare in Africa.
Limitations
: A limitation of our work is the restriction of the population surveyed to one that possessed advanced English comprehension with a relatively high degree of technical literacy and education. We believed this to be necessary in order to engage most effectively with the survey. However, we acknowledge that this may limit the extent to which the findings of this study generalize. As such, it is important that future work extends our approach with more accessible methods such as on-the-ground verbal interactions that seek to understand the perspectives of those not well-represented in our study, and extend it to local languages. Furthermore, as our study primarily leverages qualitative insights from interviews and surveys, an important area of future research involves integration of these findings with real-world evidence of the relationship between specific axes of disparities and AI-enabled health applications.
6. Conclusion
7. Reflexitivity
Two researchers work external to the researching organization, but have a combined extensive background in bioethics, philosophy, and AI for social good in Africa. Five of the other researchers work internally at the organization in various departments. One has an academic background in Sociology and Cultural Studies, one has a background in biomedical engineering, AI and global health, three have backgrounds in machine learning, statistics and algorithmic fairness in machine learning. We also varied in age, race, national origin, and sexual orientation which influenced our experiences with, and perceptions of, health and AI. Two of the co-first authors are of African origin, as is one of the additional authors. Three of the study’s authors are of South Asian and/or Middle Eastern origin. Three of the study’s authors are of North American origin.
Acknowledgements.
The authors would like to acknowledge the insights and feedback from Maya Leventer-Roberts, Sunny Jansen, Chirag Nagpal and Vishwali Mhasawade. We would also like to thank Gutcheck for their rigor, timeliness and reach in administering and analyzing the general population surveys, and D’Well for their transcription services. This study was funded by Google LLC and/or subsidiary thereof (Google). All the authors are or were employees of Google and may own stock as a part of a standard compensation package.
Appendix
Appendix A.1 General population quantitative survey results
A.1.1. AI definitons and familiarity
Figure A.1. General population responses to likert scale questions on agreement with AI definitions and familiarity with AI. a) Responses to questions poised in a likert scale format b) Responses to questions poised in a semantic differential scales format.
Agreement with provided definitions:
We found that the Gen Pop. respondents generally agreed (81-96%) with the standard definitions provided for AI, Machine Learning, Fairness and bias (Fig. A.1). Please see Appendix A.3.1 for the specific definitions provided. Of the provided definitions, bias in AI had the highest levels of strongly disagree, disagree and neutral.
Familiarity with AI and use cases:
Most of the Gen Pop. respondents were familiar with AI and its associated use cases with 11% indicating they were familiar with AI and use it regularly, 43% indicating that they had read a lot about AI and used it themselves, 31% indicating they were moderately familiar with AI but have not read much about it or used it themselves, 14% indicating they had heard of AI conceptually but had not used it, and only 2% indicating that they were not familiar at all (Fig.A.1b). We purposefully screened for respondents who had some familiarity of AI and could therefore provide knowledgeable, informed answers/opinions to the follow on questions. However this also biases our results against the section of the population which may be less educated and less aware of AI.
A.1.2. Types of questions respondents would ask an AI Chatbot:
Respondents were asked to optionally list 3 examples of questions they would ask an AI chatbot. This generated 520 questions which were analyzed and grouped into categories in a word cloud (Fig. A.2). The top categories around which questions fell into were 1) achieving health goals, e.g., “…I want to start a fitness routine. What exercises and workout schedule would be most effective for achieving my specific health goals.” [Rwanda, 30-39, Male]; (2) causes of diseases, e.g., “A sharp pain in the upper left side of the body. What are the possible causes and solutions?” [Ghana, 30-29]; and (3) symptoms of diseases, e.g., “What are the symptoms of a certain medical condition?” [Kenya, 21-29].
Figure A.2. Word cloud distribution of question types that Gen Pop. respondents would ask an AI chatbot, with examples of questions submitted. Larger size indicates high occurrence of questions type.
A.1.3. Comparisons of expert and general population survey on AI axes of disparities